### v1

# MatCLIP: Light- and Shape-Insensitive Assignment of PBR Material Models

Abstract.
Assigning realistic materials to 3D models remains a significant challenge in computer graphics. We propose MatCLIP, a novel method that extracts shape- and lighting-insensitive descriptors of Physically Based Rendering (PBR) materials to assign plausible textures to 3D objects based on images, such as the output of Latent Diffusion Models (LDMs) or photographs. Matching PBR materials to static images is challenging because the PBR representation captures the dynamic appearance of materials under varying viewing angles, shapes, and lighting conditions. By extending an Alpha-CLIP-based model on material renderings across diverse shapes and lighting, and encoding multiple viewing conditions for PBR materials, our approach generates descriptors that bridge the domains of PBR representations with photographs or renderings, including LDM outputs. This enables consistent material assignments without requiring explicit knowledge of material relationships between different parts of an object. MatCLIP achieves a top-1 classification accuracy of 76.6%, outperforming state-of-the-art methods such as PhotoShape and MatAtlas by over 15 percentage points on publicly available datasets. Our method can be used to construct material assignments for 3D shape datasets such as ShapeNet, 3DCoMPaT++, and Objaverse. All code and data will be released.
PBR Materials, Alpha-CLIP, Latent Diffusion, Transformer, Material Descriptor
copyright: nonejournal: TOGccs: Information systems Information retrievalccs: Computing methodologies Machine learningccs: Computing methodologies Texturing
1. Introduction
Creating realistic 3D models involves shaping the geometry and assigning appropriate materials and textures. While many collections of high-quality 3D shapes are available, the challenge of assigning materials that enhance realism and coherence persists. Current methods for material assignments, such as automatically selecting from a material library, often result in unsatisfactory and incoherent outcomes. While these procedural methods ensure material assignments adhere to predefined substance categories, they lack the ability to consider interdependencies between parts or align assignments with geometric and stylistic properties (cf. Fig. 1, right). Our method addresses these limitations by leveraging diffusion models and robust embeddings to produce coherent, contextually appropriate material assignments. Another approach involves matching 3D shapes to stock photos with similar textures, which can result in geometric mismatches and incorrect material assignments. Additionally, one is limited to the pre-existing designs in the stock photos. Generative models allow the exploration of a broader and less constrained space of images, a feature we aim to exploit.
Conditional Latent Diffusion Models (LDMs) (Zhang et al., 2023; Rombach et al., 2021), such as Stable Diffusion and Flux, have shown promise in generating plausible RGB images from a depth image of a 3D shape and a textual prompt describing the desired appearance and materials. However, these models produce a static image and not a 3D model. They cannot directly generate texture-based materials, such as Physically Based Rendering (PBR) materials often used in computer graphics, to produce high-quality models suitable for real-time rendering.
Figure 2. (left) Our method learns robust material descriptors through diverse renderings across geometric shapes and lighting conditions. (bottom center) MatCLIP aligns material embeddings with corresponding parts of 3D objects using cosine similarity and attention mechanisms. (right) We train our MatCLIP model using a large dataset of masked renderings. (top center) We leverage the output from LDMs (e.g., Stable Diffusion) to obtain meaningful material assignments to untextured shape collections.
To address these limitations, we propose a novel material classification model by extending Alpha-CLIP. Assigning materials from an existing material database has several advantages. First, several sources for high-quality and high-resolution texture-based PBR materials are available. For our material database, we leverage MatSynth (Vecchio and Deschaintre, 2024), a publicly available resource containing over 4,000 high-resolution (4K) PBR materials across diverse categories, including maps for basecolor, diffuse, normal, height, roughness, metallic, specular, and opacity. Second, materials from diffusion-based PBR texture generators like MatFuse (Vecchio et al., 2024) are usually unable to generate textures at such high resolutions, or they struggle with generating maps with a similar level of geometric detail.
Our contributions are as follows:
  1. (1)
    We introduce a novel method for learning embeddings of PBR materials that are robust to shape and lighting. These embeddings, trained on diverse renderings, enable robust and consistent material assignment across varied geometries and lighting conditions.
  2. (2)
    We provide an open and reproducible framework that relies on public datasets (MatSynth and 3DCoMPaT++) and foundational diffusion models, ensuring accessibility and transparency for the community.
  3. (3)
    We achieve state-of-the-art performance in material assignment, with a Top-1 accuracy of 76.69% on MatSynth, surpassing prior methods such as PhotoShape and MatAtlas.
2. Related Work
Creating realistic materials for 3D models is a longstanding challenge in computer graphics, addressed through diverse approaches such as exemplar-based material assignment, RGB-only texturing, SVBRDF estimation, and diffusion-based material generation. Exemplar-based methods like PhotoShape (Park et al., 2018) classify material categories using photographs but lack robust embeddings for PBR materials, limiting generalization. Extensions like PhotoScene (Yeh et al., 2022) optimize material parameters through differentiable graphs but depend on proprietary Adobe Substances, reducing reproducibility. In contrast, our method uses contrastive learning to align embeddings across diverse lighting and geometric conditions, improving robustness and accessibility.
Recent texture generation techniques, such as Text2Tex (Chen et al., 2023), TEXTure (Richardson et al., 2023), and MatAtlas (Ceylan et al., 2024), leverage depth-conditioned diffusion models to generate RGB textures, but face challenges with consistency across parts and geometric mismatches. Similarly, SVBRDF estimation methods, including MaterialGAN (Guo et al., 2020) and SurfaceNet (Vecchio et al., 2021), infer reflectance parameters but require complex architectures and extensive training. Our method sidesteps these challenges by directly assigning materials from a curated database, avoiding the computational overhead of texture generation and reflectance estimation.
Generative approaches like ControlMat (Vecchio et al., 2023) and MatFuse (Vecchio et al., 2024) generate new textures, with ControlMat effectively addressing high-resolution outputs but lacking emphasis on geometric accuracy, while MatFuse struggles with resolution and consistency for complex geometries. SAMa (Fischer et al., 2024) and MaterialPicker (Ma et al., 2024) focus on video-based diffusion models for material assignments. FlashTex (Deng et al., 2024) and PaintIt (Youwang et al., 2024) introduce additional complexity by requiring multi-view inputs or differentiable rendering. Unlike these methods, we emphasize reproducibility and practicality by leveraging publicly available datasets like MatSynth (Vecchio and Deschaintre, 2024) and 3DCoMPaT++ (Li et al., 2022; Slim et al., 2023), a shape dataset with over 10 million stylized 3D shapes across 41 categories.
StyleTex (Xie et al., 2024) uses diffusion models with content-negative prompts and ControlNet for stylized texture generation, focusing on aesthetics over physically-based materials. TexOct (Liu et al., 2024) synthesizes textures directly in 3D space using octree representations, addressing UV mapping and occlusion but lacking the simplicity and realism of pre-existing high-resolution PBR materials, as used in our approach. Meta 3D TextureGen (Bensadoun et al., 2024) employs a two-step text-to-image and UV refinement pipeline for globally consistent textures, but it does not leverage pre-existing high-resolution PBR materials, as our method does, ensuring consistency and realism without iterative refinement. TEXGen (Huo et al., 2025) incorporates hybrid 2D-3D denoising for UV texture synthesis, addressing geometric inconsistencies but retaining the computational overhead of UV mapping, which our method avoids by directly aligning materials with 3D shape parts. FlexiTex (Jiang et al., 2024) integrates normal maps into diffusion models for geometry-aware texture alignment but requires model-specific fine-tuning, whereas our method achieves generalization and consistency across diverse shapes using robust, pre-trained embeddings and publicly available datasets.
Our approach provides a scalable solution for assigning PBR materials, combining robust embeddings with shape- and lighting-invariant representations. By addressing limitations in material classification and ensuring coherence across diverse conditions, we position MatCLIP as a practical alternative to state-of-the-art methods for real-world applications in computer graphics.
3. Method
We begin by introducing the database of PBR materials used in this study. Unlike prior work, we rely on publicly available data to ensure repeatability and accessibility. This limits the variety of PBR materials but ensures the approach remains applicable to widely accessible resources. Next, we describe the representation of materials. We employ two different types of representations. Each shape part is modeled as a single CLIP-like vector, while each material is represented as a set of vectors capturing its appearance variations under different lighting and surface geometry. We then discuss the training process. It uses two encoders: one processes images of masked parts within multipart textured objects (Fig. 2, right); the other handles the diverse representations of PBR materials rendered under varying conditions (Fig. 2, left). Finally, we leverage pretrained diffusion models to guide material selection, utilizing their latent understanding of interdependent material assignments across parts and constraining the process to align with the specific geometry and stylistic features of the 3D model.
3.1. Material Database
A PBR material captures a diverse set of potential appearances, requiring a robust representation that is less sensitive to lighting and shape. To achieve this, we train using samples rendered under diverse conditions.
We use $n_{\text{env}}=7$ HDR environment maps from Poly Haven to represent different lighting setups for studio scenarios. Since our study focuses on furniture (chairs), we select HDR images that resemble typical indoor lighting. We further choose $n_{\text{shapes}}=6$ base shapes, including cones, cylinders, planes, spheres, Suzanne (an ape model), and tori, to ensure that material representations are not overly sensitive to shape variations. This results in $n_{\text{env}}\times n_{\text{shapes}}=42$ renderings per material. For chairs, MatSynth includes 5 applicable substance categories, such as fabric and leather, which span a total of 1,600 materials. By rendering each material across all 42 lighting and shape combinations, we generate 67,200 images, capturing various lighting and geometric conditions while maintaining computational efficiency.
Using the MatSynth material database, we render each material with Blender’s physically-based path tracer. Each shape uses a fixed UV-mapping. This choice is practical, as rendering is computationally expensive. Additionally, consistent UV-mappings ensure physical scale across materials, avoiding the need to make the model robust to varying UV-mappings, which would require significant additional complexity.
We render each shape from a frontal view using a focal length of $100mm$, chosen to match the focal lengths commonly used in professional product photography, such as IKEA catalogs. The distance is set to maximize the shape’s extent within the image margins, ensuring that material details are clearly visible. Figure 2 (left) shows a subset of renderings for one material applied to various shapes under various environment maps.
We restrict our evaluation to chairs in the 3DCoMPaT++ (Slim et al., 2023) dataset, but the pipeline is suitable for all furniture and other products used in daily life. Chairs are well-suited for studying material assignments due to their geometric diversity and the presence of distinct material groups. This focus aligns with prior work, such as PhotoShape (Park et al., 2018), which also concentrates on chairs.
Training Data
To generate the dataset (cf. Fig. 2 (right)), we first randomly select a shape, then assign random materials to its parts based on their annotated substances. Each shape may appear multiple times with different material configurations. Images are rendered from an informative perspective, selected randomly from a range of viewpoints so that key features are likely to be visible. We apply random rotations, sampling typical viewing angles with $360^{\circ}$ azimuth and $0^{\circ}$$90^{\circ}$ elevation (no roll). Models are rendered under the same set of HDR environment maps used in the material database. For each image, we also render a mask for each material group defined in the 3DCoMPaT++ model. This process yields a dataset of 53,000 rendered images. However, each image contains multiple material groups, with masks indicating the parts where specific materials are applied. The total dataset includes over 200,000 individual parts (each a material group within an image), which serve as training samples.
Figure 3. Examples (in each row from left to right) of 3D models, masks, depth images, edge images, and corresponding Stable Diffusion outputs.
3.2. Material Representation
To represent materials in our database, we require embeddings that are less sensitive to lighting, shape, and background context. While Alpha-CLIP (Sun et al., 2023) provides an effective mechanism for masking regions of interest in static images, it is unsuitable for representing the materials themselves. Alpha-CLIP embeddings retain information about the mask’s shape and the unmasked context, which can lead to undesirable incorporation of irrelevant background features when applied to material renderings.
Instead, we adopt a different preprocessing strategy for material representations. Since our material renderings use well-defined reference shapes (e.g., spheres, cones, planes, tori), we crop the rendered image to isolate only the textured region, ensuring that the embedding focuses entirely on the material. Specifically, we compute the largest inscribed rectangle within the region of interest, ensuring that the crop contains only material textures without any background pixels. The cropped region is then passed through a standard CLIP image encoder to generate a descriptor for each rendering (Fig. 2, left).
To account for the variation introduced by lighting conditions, shading, and texture distortions across different shapes, we compute CLIP embeddings for the full set of $n_{\text{env}}\times n_{\text{shapes}}=42$ rendered images per material. These embeddings are arranged into a sequence and input to the material representation branch of our model (Fig. 2, left), allowing it to learn a distribution of appearances that is more robust to diverse conditions.
Method Top-1 Accuracy [%] $\uparrow$ Top-5 Accuracy [%] $\uparrow$
PhotoShape (ResNet-34) 57.21 77.39
PhotoShape (ResNet-50) 58.76 77.62
PhotoShape (ViT-B/16) 60.86 78.23
OpenCLIP (Max. Inscribed Square Crop, Max. Score to 42 Material Features) 16.13 31.49
OpenCLIP (Max. Inscribed Square Crop, Avg. Material Features) 19.90 35.34
MatAtlas (0.8 × LAB + 0.2 × CLIP, Max. Score to 42 Material Features) 18.77 36.79
MatAtlas (0.8 × LAB + 0.2 × CLIP, Avg. Material Features) 13.14 28.40
MatCLIP (Ours) 76.69 89.02
Table 1. Quantitative evaluation of our method. We compare to variants of PhotoShape (Park et al., 2018), simple crop matching using CLIP (Radford et al., 2021), and variants of the core method of MatAtlas (Ceylan et al., 2024). We either average or identify the max. score across the 42 material features of each of our materials in case of CLIP and MatAtlas. Accuracy is evaluated on Top-1 and Top-5 metrics, with higher values indicating better performance.
3.3. MatCLIP Architecture
Our MatCLIP model consists of two main encoders, as illustrated in Fig. 2 (bottom center): the Material Encoder, denoted $E_{\text{material}}$, and the Part Encoder, denoted $E_{\text{part}}$. The Material Encoder processes a set of rendered images of a PBR material, while the Part Encoder processes a masked region of an input image. The goal of MatCLIP is to align materials, represented by their diverse appearances under varying conditions, with the corresponding masked parts in input images.
The architecture is inspired by Alpha-CLIP (Sun et al., 2023), which aligns text with masked portions of an image. In MatCLIP, we replace Alpha-CLIP’s text-processing encoder with $E_{\text{material}}$, which processes sets of rendered PBR images. The embeddings generated by $E_{\text{material}}$ leverage the 42 renderings of each material, as described in Section 3.2. These embeddings are passed through a transformer to produce a robust representation of the material. In parallel, $E_{\text{part}}$ processes masked regions of input images, outputting embeddings that represent the appearance of the part in its context.
During training, embeddings from $E_{\text{material}}$ and $E_{\text{part}}$ are aligned using a contrastive loss, following the approach used in CLIP (Radford et al., 2021). This loss encourages the model to associate corresponding embeddings while distinguishing unrelated pairs.
3.4. Sampling Textures from Stable Diffusion
Generating plausible texture assignments to 3D objects is challenging because we do not know the true distribution of material assignments. LDMs offer a powerful approach to assigning plausible materials to 3D models by capturing realistic texture distributions from large-scale datasets. This enables material assignments that are both coherent and diverse.
In our approach, a target image, denoted $I_{\text{target}}$, is a synthetic rendering produced by Stable Diffusion (Rombach et al., 2021; Esser et al., 2024) (or Flux (Labs, 2023)) that combines geometric information from a 3D model with a stylistic interpretation of its materials. The purpose of $I_{\text{target}}$ is to represent a visually plausible appearance for the 3D model that aligns with its structure, as shown in Fig. 3 (right).
To generate $I_{\text{target}}$, we condition Stable Diffusion on depth and edge information derived from the 3D model, ensuring that the generated textures align with the model’s geometry. Depth information is captured by rendering a single-view depth image, $I_{\text{depth}}$, from an informative perspective, as discussed in Section 3.2. In addition to depth conditioning, edge information is extracted using Canny edge detection on a diffuse gray rendering of the 3D model. This ensures fine geometric details, such as thin wires or fabric strips, are preserved in $I_{\text{target}}$. We also incorporate a CLIP text embedding derived from descriptive prompts suggested by ChatGPT (e.g., “a sturdy and durable chair with a robust design”), which guides the stylistic interpretation of the material. By combining depth, edge, and text conditioning, Stable Diffusion produces target images that capture both geometric accuracy and material diversity. Examples of depth images, edge maps, and resulting outputs are shown in Fig. 3.
Once $I_{\text{target}}$ is generated, we create part-level masks to isolate regions for texture assignments. These masks enable precise alignment between generated textures and specific model regions. Each part is assigned a unique mask, as illustrated in Fig. 3. Given a target image and masks, we can now use our MatCLIP model to match materials from the MatSynth dataset. The whole procedure of using Stable Diffusion together with our model is shown in Fig. 2 (top center).
4. Results
We evaluate our method across three key dimensions: (1) material classification performance, demonstrating MatCLIP’s effectiveness in identifying materials across diverse shapes and lighting conditions; (2) generalization to new inputs, including manually annotated evaluations of substances derived from latent diffusion model outputs; and (3) ablation studies assessing the impact of dataset diversity on classification accuracy. Complementary qualitative evaluations (Section 4.4) illustrate MatCLIP’s ability to extend beyond known material distributions. Together, these experiments validate MatCLIP’s robustness in procedural modeling tasks and its superiority over state-of-the-art methods.
Our model is trained on a GPU cluster node with 4 NVIDIA A100 40GB cards for approximately 48 hours. Hyperparameters follow the contrastive loss strategy introduced in Alpha-CLIP (Sun et al., 2023). MatCLIP adapts this approach to align part-level embeddings with diverse material representations, as detailed in Section 3.2. This enables MatCLIP to robustly handle variations in shape and lighting conditions while maintaining consistent performance.
For quantitative evaluations, we partition the dataset at the object level to ensure no object appears in both training and test sets, even when rendered with different material assignments. This design prevents model overfitting to specific shapes and ensures unbiased evaluations. While all materials are included in both training and test sets, the object-level separation highlights MatCLIP’s ability to generalize across new shape configurations within a closed material set.
Qualitative evaluations presented in Section 4.4 further demonstrate MatCLIP’s ability to generalize to novel material distributions. These results highlight MatCLIP’s potential to extend beyond quantitative metrics, reinforcing its applicability to real-world procedural modeling tasks.
4.1. Classification Performance
We evaluate MatCLIP’s material classification performance by ranking materials in the test split of our dataset, using descriptors generated by the part encoder ($E_{\text{part}}$) and material encoder ($E_{\text{material}}$). Materials are ranked by cosine similarity between their embeddings. For each query shape from the test set, materials were matched from the entire set of available materials, without restricting classification to the substance categories defined for the 3DCoMPaT++ shapes. This ensured that the model was evaluated across the full database of materials. The Top-1 and Top-5 accuracies of our method, compared to PhotoShape (Park et al., 2018) and MatAtlas (Ceylan et al., 2024), are shown in Table 1.
For a fair comparison, we modified PhotoShape by replacing its original ResNet34 and ResNet50 backbones with the more recent ViT-B/16 architecture to better align with modern methods. This resulted in slightly improved performance of PhotoShape on our data. Additionally, we evaluated MatAtlas without restricting classification to the material classes suggested by ChatGPT, instead testing its ability to match materials from the full database. Although this adjustment reduced MatAtlas’s performance, our intent was to assess only the classification component, ensuring the comparison focused on embedding quality rather than substance category limitations. These adjustments aimed to standardize the evaluation setup and ensure all methods were tested under similar conditions. Despite these efforts to strengthen the baselines, MatCLIP achieves the highest accuracy, with a Top-1 accuracy of 76.69% and a Top-5 accuracy of 89.02%, significantly outperforming the next-best method, PhotoShape (ViT-B/16) which had a top-1 accuracy of 60.86%.
These results underscore the robustness of $E_{\text{part}}$ and $E_{\text{material}}$ in generating descriptors that effectively handle the complexity of diverse objects and material assignments, establishing MatCLIP as a state-of-the-art approach for material classification.
4.2. Generalization to Unseen Materials
To evaluate MatCLIP’s performance in assigning materials not explicitly part of the training set, we implemented a GUI for manual annotation of outputs from Stable Diffusion. This was needed because the ground truth was not known for LDM outputs. It was used to assign the most likely substance to highlighted shape parts. In total, approximately 400 annotations were collected, covering various parts from 100 different 3DCoMPaT++ chair models.
Method Val.Acc. [%] $\uparrow$
PhotoShape (ResNet-34) 46.67
PhotoShape (ResNet-50) 49.07
PhotoShape (ViT-B/16) 44.27
OpenCLIP (Max. Score) 49.20
OpenCLIP (Avg. Features) 44.65
MatAtlas (0.8 × LAB + 0.2 × CLIP, Max.) 52.94
MatAtlas (0.8 × LAB + 0.2 × CLIP, Avg.) 48.93
MatCLIP (Ours) 55.47
Table 2. Model comparison using manually annotated ground-truth data. We evaluate the performance of assigning the correct substance derived from matched materials. We use the same models as in Table 1.
We then compared MatCLIP’s Top-1 ranked material assignments to those of competing methods, including PhotoShape (Park et al., 2018) and MatAtlas (Ceylan et al., 2024). The results are summarized in Table 2. Although MatCLIP consistently outperforms other approaches, the performance gap is narrower than in prior evaluations. This is likely due to the inherent ambiguity and subjectivity in manual annotations.
For a fair comparison, the same baseline configurations were used as in Section 4.1. Notably, we evaluated each method’s ability to rank substances from the full material database, without restricting classification to substance categories predefined in 3DCoMPaT++. Despite this broader evaluation scope, MatCLIP achieves a Top-1 accuracy of 55.47%, outperforming all competitors.
4.3. Ablation Study
To assess the impact of dataset diversity on material classification accuracy, we conducted an ablation study by varying the number of shapes and environment maps used to render materials. The full MatCLIP model relies on 42 renderings per material, generated using 6 distinct shapes and 7 HDR environment maps. We systematically reduced the dataset by testing subsets, including cases with a single shape across all environment maps (Plane 7 Env) and all shapes under a single environment map (All Shapes 1 Env). For each subset, we retrained MatCLIP and monitored Top-1 classification accuracy over the first 9,000 training steps.
The results, shown in Table 3, highlight the significant advantage of using the full dataset. The full model achieves the highest accuracy, significantly outperforming models trained on subsets. Notably, the worst performance was observed with a single shape under a single environment map (Plane 1 Env), emphasizing the importance of both geometric and lighting diversity for learning robust material representations.
These findings confirm that diverse shapes and lighting conditions are crucial for MatCLIP’s performance. Future improvements could leverage an even broader set of base shapes and environment maps to further enhance material generalization and classification robustness.
Model Name Val. Acc. [%] $\uparrow$
Plane Shape, 1 Environment Map 7.01
Plane Shape, 7 Environment Maps 53.19
All Shapes, 1 Environment Map 56.68
MatCLIP (Ours), Full Model 60.54
Table 3. Ablation study examining the effect of shape and environment map diversity on classification accuracy. The full model, which uses all six shapes and seven environment maps (42 renderings per material), performs best. Variants include: ’Plane 1 Env,’ using a single shape (plane) and one environment map; ’Plane 7 Env,’ using the plane with all seven environment maps; and ’All Shapes 1 Env,’ using all six shapes with a single environment map.
4.4. Qualitative Evaluations with Latent Diffusion Models
We conduct qualitative evaluations using outputs from Stable Diffusion (cf. Fig. 7) and Flux (cf. Fig. 8) to assess MatCLIP’s ability to generalize beyond the closed set of training materials. These evaluations focus on assigning materials to synthetic target images ($I_{\text{target}}$) generated using depth, edge, and text conditioning, as described in Section 3.2. Our model selects the best-matching material from the MatSynth database for each target image, producing visually coherent and contextually appropriate assignments, even for complex shapes and lighting conditions.
Figures 7 and 8 showcase examples where MatCLIP robustly matches materials to parts of a 3D shape. The visual results illustrate the effectiveness of our embeddings in achieving plausible and realistic material assignments across diverse target distributions.
4.5. Prompt-Based Design Exploration
In order to explore a larger design space, we leverage the power of ChatGPT to generate a large number of prompts. In general, we distinguish between two types of prompts. We use a regular prompt that usually only consists of a few words like ”a comfortable chair.” The prompts we try to obtain from ChatGPT we refer to as advanced prompts, since they usually consist of very detailed descriptions of materials and the relationship between the materials that are used for different parts of a shape. To allow ChatGPT to describe such relationships, we do not only use a textual prompt but also provide an exemplar image of the particular shape of interest (cf. Fig. 4). Using advanced prompts, we are able to obtain a larger variety than we were able to get using only regular prompts (cf. Fig. 9).
Refer to caption This image was generated by an LDM using the prompt ”A photograph of a very comfortable chair, high quality, white background.” Please suggest 3 different prompts to obtain more meaningful but very differently looking results that take the geometry into account. Here are three alternative prompts you can use to generate varied but geometrically related chairs using a latent diffusion model: (1) A futuristic armchair inspired by mid-century modern design, featuring metallic legs and a sleek seat. (2) A luxurious lounge chair with an organic design, curved lines, and upholstered with soft velvet fabric. (3) A modern Scandinavian-style chair with a minimalist wooden frame, light fabric cushions, and angled legs.
Figure 4. We exploit ChatGPT to obtain various prompts we then use to condition Flux. Top to bottom: The query target image, our query prompt, and the response from ChatGPT representing a set of advanced prompts we use to obtain the material variations shown in Fig. 9, middle row.
4.6. User Study
To evaluate the effectiveness of our method for assigning plausible materials to large non-textured shape collections, we conducted a user study on Amazon Mechanical Turk (MTurk). Specifically, we selected 100 chair shapes at random from the 3DCoMPaT++ dataset and used Stable Diffusion, as detailed in Section 3.4, to generate plausible material assignments. For each chair shape, we created approximately five different sets of material assignments, resulting in a total of 506 textured chair shapes.
To benchmark our approach, we compared it to the procedural material assignment method from 3DCoMPaT++, which procedurally assigns materials to shape parts based on predefined substance categories. All chair shapes, including those textured using our method and the procedural approach, were rendered using Blender’s physically-based path tracer. To ensure a fair comparison, we used identical rendering parameters for all models, including the same camera settings (intrinsic and extrinsic), environment map, and overall scene composition.
We constructed a dataset of image pairs, each consisting of one chair textured using our method and one textured using the procedural approach. The order of images within each pair was randomized, and participants on MTurk were asked to choose the chair they believed had the more plausible material assignment (cf. Fig. 6). We collected two independent responses (HITs) per image pair to increase reliability, resulting in 1,012 ranked decisions.
The results demonstrate a clear preference for our method: in 783 of the 1,012 pairs (77.37%), participants selected the chair textured using our approach as having more plausible material assignments. In contrast, the procedural method from 3DCoMPaT++ was preferred in only 229 pairs (22.63%). These findings underscore the effectiveness of our approach, leveraging the material distributions suggested by Stable Diffusion, in producing more realistic and contextually appropriate textures for large shape collections compared to existing procedural methods.
5. Limitations
Despite MatCLIP’s strong performance in assigning materials to 3D shape parts, our method has certain limitations. One challenge arises when Stable Diffusion introduces geometric artifacts, such as fine-structured geometric details not present in the original shape, despite conditioning on depth and edge inputs (cf. Fig. 5). These artifacts can occasionally lead to material assignments that appear inconsistent with the input.
Generalization to unseen material distributions remains an open challenge. While qualitative evaluations (cf. Section 4.4) suggest that MatCLIP can produce visually coherent material assignments for unseen scenarios, quantitative evaluations using manually annotated substances show room for improvement. The Top-1 accuracy drops from 76.69% on the validation set of known materials (cf. Table 1) to 55.47% for manual annotations (cf. Table 2). However, this reduction is not unique to our method; other methods also experience performance drops under similar conditions. We believe this is mainly due to the uncertainty in the ground truth annotations, as the task of material assignment from RGB images is also very difficult for human labelers.
Figure 5. Our model struggles in cases where Stable Diffusion adds additional geometry (e.g., wrinkles on the left side) or where no material from the material database would give a good match (right side).
6. Conclusion
This paper introduced MatCLIP, a novel framework for robust material classification and assignment to 3D objects. Our method achieves state-of-the-art performance in material classification, leveraging embeddings that are less sensitive to variations in shape and lighting. By training on a diverse dataset of renderings, MatCLIP demonstrates the ability to handle complex material representations, achieving a Top-1 accuracy of 76.69% on MatSynth, significantly outperforming prior top methods such as PhotoShape and MatAtlas. This contribution establishes a new SOTA for robust material classification in diverse geometric and lighting conditions.
Figure 6. Layout of the user study on MTurk. Workers were asked to choose the more plausible material assignment. We compare our method to the procedural assignments from 3DCoMPaT++.
Beyond classification, we demonstrated MatCLIP’s practical application in assigning materials to 3D objects, particularly in scenarios where ground-truth materials are unknown. By integrating Latent Diffusion Models (LDMs) to generate target images, our approach assigns coherent and visually plausible materials, addressing a key challenge in procedural modeling. Qualitative evaluations show that MatCLIP can generalize to unseen material distributions, effectively bridging the gap between synthetic outputs and high-resolution PBR material databases.
By building on publicly available resources like MatSynth and 3DCoMPaT++, MatCLIP ensures accessibility and reproducibility, providing an open framework for the community.